2 Liquidity Mismatch Consider a bank whose assets are composed of risky assets and Treasuries (both with 10-year durations) and whose liabilities are made up of equity and debt. To begin with, suppose that on the asset side, the bank holds 80% risky assets and 20 Treasuries, while on the liabilities side, the bank has 90% short-term debt (deposits) and 10% equity. In the following questions, you will be asked how the banks portfolio composition changes when it takes certain actions. Use these numbers as the reference portfolio for all questions 1. Suppose that the bank sells Treasuries and uses the proceeds to buy risky assets such that after the swap, its holdings are 95% risky assets and 5% Treasuries. Assume the bank issues no additional liabilities.(a) How does the banks leverage change? b) How does the banks maturity mismatch change? (c) How does the banks liquidity mismatch change? 2. Now suppose that the bank issues additional deposits and uses the proceeds to buy Treasuries so that its portfolio is composed of 40% risky assets and 60% Treasuries (a) What is the composition of the banks liabilities now? How does the banks leverage change? (b) How does the banks maturity mismatch change? Assume the banks equity has a duration of 20 vears. (c) How does the banks liquidity mismatch change?
Answer #1

ANSWER

a)Bank sells treasuries and use this fund to buy risky assets .The bank becomes more levered and it increases its debt and reduces its equity

Bank leverage= ((90/80))*95=106.875 debt

equity= ((10/20))*5=2.5 equity.

b)maturity mismatch is a financial situation of a financial institution or company in which assets held to meet future liabilities are not aligned in terms of maturity time. How a company organizes the maturity of its assets and liabilities can give details into the liquidity of its position. When there is a material maturity mismatch, a liquidity squeeze could arise.

In the above case Debt increased from 90 to 106.875

equity decreased from 10 to 2.5

C) As bank issues no additional liability there is no liquidity mismatch.

2

a).Bank issues additional deposits and proceeds to buy treasuries.

compodition of banks liabilitu now:- ((90/80))*40=45 debt

and equity become = ((10/20))× 60=30 treasuries

still the company is high levered ad debt is more.

b) Banks maturity mismatch

bank debt decresed from 90 to 45

and equity incresed from 20 to 30

(c)bank issue additional deposits to buy treasuries.it has looses in liquidity postion since treasuries have life 20 years.